OpenStudy (anonymous):

Find the x-intercepts for the graph of the equation y=x^2-3x-4 The x-intercepts are?

OpenStudy (immanuelv):

-1 and 4

OpenStudy (amistre64):

-(-4,1) x = 4,-1

OpenStudy (amistre64):

got yor signs mixed on that one

OpenStudy (anonymous):

so when writing it the 4 comes first then the -1 right?

OpenStudy (amistre64):

doesnt matter

OpenStudy (immanuelv):

sorryy (X+1)(X-4)

OpenStudy (amistre64):

4,-1 -1,4 x=x

OpenStudy (anonymous):

Find the x-intercepts for the graph of the equation y=x^2-3x-4 The x-intercepts are? just equate the equation to o, or y=0

OpenStudy (amistre64):

its like saying which shoe do you put on first....

OpenStudy (immanuelv):

left one jk

OpenStudy (anonymous):

gotcha thanks guys

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

;)

OpenStudy (anonymous):

y=x^2-3x-4 0=x^2-3x-4 (x-4)(x+1)=0 x=4 x=-1

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

those are the x intercepts.

OpenStudy (anonymous):

you are welcome.